# Tensor Product Attention Is All You Need

### v1

@@ 1 Introduction
Figure 1: Tensor Product Attention (TPA) in the Tensor ProducT ATTenTion Transformer (T6). Different from multi-head attention, in each layer, firstly the hidden state goes through different linear layers to get the latent factor matrices $\mathbf{A}$’s and $\mathbf{B}$’s for query, key, and value. We additionally apply RoPE to $\mathbf{B}_{Q}$ and $\mathbf{B}_{K}$ for query and key. Then the multi-head query, key, and value vectors are attained by the tensor product of $\mathbf{A}_{(\cdot)}$ and $\mathbf{B}_{(\cdot)}$. Finally, the output of TPA is produced by scaled dot-product attention followed by linear projection of concatenated results of multiple heads.
A variety of solutions have been explored to address this memory bottleneck. Some approaches compress or selectively prune cached states through sparse attention patterns (Child et al., 2019) or token eviction strategies (Zhang et al., 2023c; Xiao et al., 2024; Ribar et al., 2024), though such methods risk discarding tokens that may later prove important. Other work proposes off-chip storage of key-value states (He & Zhai, 2024), at the expense of increased I/O latency. Attention variants like multi-query attention (MQA) (Shazeer, 2019) and grouped-query attention (GQA) (Ainslie et al., 2023) reduce per-token cache requirements by sharing keys and values across heads, but often compromise flexibility or require significant architectural modifications. Meanwhile, low-rank weight factorization methods such as LoRA (Hu et al., 2022) effectively reduce fine-tuning memory, yet do not address the KV cache overhead that dominates runtime. The recently introduced Multi-head Latent Attention (MLA) in Deepseek-V2 (Liu et al., 2024a) caches compressed key-value representations but needs additional position-encoded parameters per head due to incompatibility with Rotary Position Embedding (RoPE) efficiently (Su et al., 2024b).
@@
A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (Su et al., 2024b), enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (Touvron et al., 2023) and Gemma (Team et al., 2024).
@@
  • We propose Tensor ProducT ATTenTion Transformer (T6), a new TPA-based model architecture for sequence modeling. On language modeling experiments, T6 consistently improves validation perplexity and downstream evaluation performance with reduced KV cache size.
  • We show TPA integrates seamlessly with RoPE (Su et al., 2024b), facilitating easy adoption in popular foundation model architectures such as LLaMA and Gemma.
(a) Training Loss
(b) Validation Loss
Figure 2: Training loss and validation loss of pretraining large-size (773M) models with different attention mechanisms on the FineWeb-Edu-100B dataset.
@@ 2 Background
In this section, we review several classical forms of attention: Scaled Dot-Product Attention, Multi-Head Attention (MHA) (Vaswani et al., 2017), Multi-Query Attention (MQA) (Shazeer, 2019), and Grouped Query Attention (GQA) (Ainslie et al., 2023), as well as Rotary Position Embedding (RoPE, Su et al. (2024b)). We also introduce a recent method called Multi-head Latent Attention (MLA) used in DeepSeek-V2 (Liu et al., 2024a) and DeepSeek-V3 (Liu et al., 2024b).
@@ 2.2 Multi-Head Attention (MHA)
$\mathbf{Q}_{t,i}=(\bm{W}_{i}^{Q})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},\quad\mathbf{K}_{t,i}=(\bm{W}_{i}^{K})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},\quad\mathbf{V}_{t,i}=(\bm{W}_{i}^{V})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$
@@
2.3 Multi-Query Attention (MQA)
Hence, each head $i$ only has a distinct query $\mathbf{Q}_{i}\in\mathbb{R}^{T\times d_{k}}$, but shares the same key $\mathbf{K}_{\text{shared}}\in\mathbb{R}^{T\times d_{k}}$ and value $\mathbf{V}_{\text{shared}}\in\mathbb{R}^{T\times d_{k}}$. In practice, this means:
$\bm{W}^{Q}_{i}\in\mathbb{R}^{d_{\text{model}}\times d_{k}},\quad\bm{W}^{K}_{\text{shared}},\bm{W}^{V}_{\text{shared}}\;\in\;\mathbb{R}^{\,d_{\text{model}}\times d_{k}}.$
The resulting MQA operation is:
where
2.4 Grouped Query Attention (GQA)
and
Again, $\bm{W}^{K}_{g},\bm{W}^{V}_{g}\in\mathbb{R}^{d_{\text{model}}\times d_{k}}$ for each group $g$, and $\bm{W}^{Q}_{i}\in\mathbb{R}^{\,d_{\text{model}}\times d_{k}}$ for each head $i$. The complete output is again a concatenation of all heads:
@@ 2.5 Rotary Position Embedding (RoPE)
Many recent LLMs use rotary position embedding (RoPE; Su et al., 2024b) to encode positional information in the query/key vectors. Specifically, let $\mathrm{RoPE}_{t}$ denote the rotation operator $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ corresponding to the $t$-th position. $\mathbf{T}_{t}$ is a block-diagonal matrix, which consists of block-diagonal matrix$\begin{pmatrix}\cos(t\theta_{j})&-\sin(t\theta_{j})\\ \sin(t\theta_{j})&\cos(t\theta_{j})\end{pmatrix}$, $j\in\{1,\cdots,d_{h}/2\}$, where $\{\theta_{j}\}$ are pre-defined frequency parameters, e.g., $\theta_{j}=1/10000^{2j/d_{h}}$. Then we define
@@
2.6 Multi-head Latent Attention (MLA)
$\mathbf{C}^{KV}$ $=\mathbf{X}\bm{W}^{DKV},\quad(\bm{W}^{DKV}\in\mathbb{R}^{\,d_{\text{model}}\times d_{c}}),$
$\operatorname{Concat}\bigl{(}\mathbf{K}_{1}^{C},\mathbf{K}_{2}^{C},\ldots,\mathbf{K}_{h}^{C}\bigr{)}=\mathbf{K}^{C}$ $=\mathbf{C}^{KV}\bm{W}^{UK},\quad(\bm{W}^{UK}\in\mathbb{R}^{d_{c}\times d_{h}h}),$
$\mathbf{K}^{R}$ $=\operatorname{RoPE}\bigl{(}\mathbf{X}\bm{W}^{KR}\bigr{)},\quad(\bm{W}^{KR}\in\mathbb{R}^{d_{\text{model}}\times d_{h}^{R}}),$
$\operatorname{Concat}\bigl{(}\mathbf{V}_{1}^{C},\mathbf{V}_{2}^{C},\ldots,\mathbf{V}_{h}^{C}\bigr{)}=\mathbf{V}^{C}$ $=\mathbf{C}^{KV}\bm{W}^{UV},\quad(\bm{W}^{UV}\in\mathbb{R}^{d_{c}\times d_{h}h}),$
where $\mathbf{C}^{KV}\in\mathbb{R}^{T\times d_{c}}$ is the compressed KV latent (with $d_{c}\ll d_{h}h$), and $\operatorname{RoPE}(\cdot)$ represents the RoPE transform applied to the separate key embeddings $\mathbf{K}^{R}$ of dimension $d_{h}^{R}$. Thus, only $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ need to be cached, reducing KV memory usage while largely preserving performance compared to standard MHA (Vaswani et al., 2017).
$\mathbf{C}^{Q}$ $=\mathbf{X}\bm{W}^{DQ},\quad(\bm{W}^{DQ}\in\mathbb{R}^{\,d_{\text{model}}\times d_{c}^{\prime}}),$
$\operatorname{Concat}\bigl{(}\mathbf{Q}_{1}^{C},\mathbf{Q}_{2}^{C},\,\ldots,\mathbf{Q}_{h}^{C}\bigr{)}=\mathbf{Q}^{C}$ $=\mathbf{C}^{Q}\bm{W}^{UQ},\quad(\bm{W}^{UQ}\in\mathbb{R}^{\,d_{c}^{\prime}\times d_{h}h}),$
$\operatorname{Concat}\bigl{(}\mathbf{Q}_{1}^{R},\,\mathbf{Q}_{2}^{R},\,\ldots,\,\mathbf{Q}_{h}^{R}\bigr{)}=\mathbf{Q}^{R}$ $=\operatorname{RoPE}\bigl{(}\mathbf{C}^{Q}\bm{W}^{QR}\bigr{)},\quad(\bm{W}^{QR}\in\mathbb{R}^{d_{c}^{\prime}\times d_{h}^{R}h}),$
Here, $\mathbf{C}^{Q}\in\mathbb{R}^{T\times d_{c}^{\prime}}$ (with $d_{c}^{\prime}\ll d_{h}h$) is the compressed query latent. As above, each $\bm{W}^{DQ}$, $\bm{W}^{UQ}$, and $\bm{W}^{QR}$ connects these lower-dimensional query latents back to $h$ heads of dimension $d_{h}+d_{h}^{R}$.
$\mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}=[(\bm{W}^{UQ}_{i})^{\top}(\bm{W}^{DQ}_{i})^{\top}\mathbf{x}_{t}]^{\top}[(\bm{W}^{UK}_{i})^{\top}\mathbf{c}^{KV}_{s}]=\mathbf{x}_{t}^{\top}[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]\mathbf{c}^{KV}_{s},$ (2.2)
[…]
Different from (2.2), acceleration by pre-computing $[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{T}_{t-s}}(\bm{W}^{UK}_{i})^{\top}]$ fails since it varies for different $(t,s)$ position pairs. Therefore, MLA adds the additional $\mathbf{k}_{t}^{R}$ part with a relatively smaller size for RoPE compatibility. In Section 3.2, we will show that TPA addresses the issue of RoPE-incompatibility by applying tensor product.
@@ 3.1 Tensor Factorization of Queries, Keys, and Values
$\mathbf{Q}_{t}$ $=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\quad$ $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h},\;\;$ $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}},$ (3.1)
$\mathbf{K}_{t}$ $=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\quad$ $\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h},\;\;$ $\mathbf{b}^{K}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}},$ (3.2)
$\mathbf{V}_{t}$ $=\frac{1}{R_{V}}\sum_{r=1}^{R_{V}}\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{V}_{r}(\mathbf{x}_{t}),\quad$ $\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h},\;\;$ $\mathbf{b}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}}.$ (3.3)
Hence, for queries, each tensor product $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\colon\mathbb{R}^{h}\times\mathbb{R}^{d_{h}}\to\mathbb{R}^{h\times d_{h}}$ adds up to form the query slice $\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}$. Similarly, analogous definitions apply to key slice $\mathbf{K}_{t}$ and value slice $\mathbf{V}_{t}$.
@@
$\mathbf{a}^{Q}(\mathbf{x}_{t})=\bm{W}^{a^{Q}}\,\mathbf{x}_{t}\in\mathbb{R}^{R_{q}\cdot h},\quad\mathbf{b}^{Q}(\mathbf{x}_{t})=\bm{W}^{b^{Q}}\,\mathbf{x}_{t}\in\mathbb{R}^{R_{q}\cdot d_{h}},$
@@ 3.2 RoPE Compatibility and Acceleration
$\widetilde{\mathbf{K}}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{(r)}(\mathbf{x}_{t})\otimes\operatorname{RoPE}_{t}\bigl{(}\mathbf{b}^{K}_{(s)}(\mathbf{x}_{t})\bigr{)}=\frac{1}{R_{K}}\mathbf{A}_{K}(\mathbf{x}_{t})^{\top}\operatorname{RoPE}_{t}\bigl{(}\mathbf{B}_{K}(\mathbf{x}_{t})\bigr{)}.$
@@
$\operatorname{RoPE}({\mathbf{Q}}_{t})=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\,\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t}),\qquad\text{where }\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t})=\operatorname{RoPE}_{t}\bigl{(}\mathbf{B}_{Q}(\mathbf{x}_{t})\bigr{)}.$ (3.7)
In addition, assume $\mathbf{Q}_{t}$ and $\mathbf{K}_{s}$ are factorized by TPA and then rotated by $\operatorname{RoPE}_{t},\operatorname{RoPE}_{s}$. Let $\widetilde{\mathbf{Q}}_{t}=\operatorname{RoPE}_{t}(\mathbf{Q}_{t})$ and $\widetilde{\mathbf{K}}_{s}=\operatorname{RoPE}_{s}(\mathbf{K}_{s})$. Then we have
@@
Theorem 1 indicates that TPA does not break RoPE’s relative translational property. We prove Theorem 1 in Appendix A. In short, $\operatorname{RoPE}_{t}$ acts as a block-diagonal orthogonal transform (i.e., a matrix $\mathbf{T}_{t}$) on $\mathbf{B}_{Q}(\mathbf{x}_{t})$. Consequently, $\mathbf{A}_{Q}(\mathbf{x}_{t})$ remains unchanged, while each column of $\mathbf{B}_{Q}(\mathbf{x}_{t})$ is rotated appropriately, preserving the TPA structure.

### v2

@@ 1 Introduction
Figure 1: Tensor Product Attention (TPA) in the Tensor ProducT ATTenTion Transformer (T6). Different from multi-head attention, in each layer, firstly the hidden state goes through different linear layers to get the latent factor matrices $\mathbf{A}_{(\cdot)}$’s and $\mathbf{B}_{(\cdot)}$’s for query, key, and value. We additionally apply RoPE to $\mathbf{B}_{Q}$ and $\mathbf{B}_{K}$ for query and key. Then the multi-head query, key, and value vectors are attained by the tensor product of $\mathbf{A}_{(\cdot)}$ and $\mathbf{B}_{(\cdot)}$. Finally, the output of TPA is produced by scaled dot-product attention followed by linear projection of concatenated results of multiple heads.
A variety of solutions have been explored to address this memory bottleneck. Some approaches compress or selectively prune cached states through sparse attention patterns (Child et al., 2019) or token eviction strategies (Zhang et al., 2023c; Xiao et al., 2024; Ribar et al., 2024), though such methods risk discarding tokens that may later prove important. Other work proposes off-chip storage of key-value states (He & Zhai, 2024), at the expense of increased I/O latency. Attention variants like multi-query attention (MQA) (Shazeer, 2019) and grouped-query attention (GQA) (Ainslie et al., 2023) reduce per-token cache requirements by sharing keys and values across heads, but often compromise flexibility or require significant architectural modifications. Meanwhile, low-rank weight factorization methods such as LoRA (Hu et al., 2022) effectively reduce fine-tuning memory, yet do not address the KV cache overhead that dominates runtime. The recently introduced Multi-head Latent Attention (MLA) in Deepseek-V2 (Liu et al., 2024a) caches compressed key-value representations but needs additional position-encoded parameters per head due to incompatibility with Rotary Position Embedding (RoPE) efficiently (Su et al., 2024).
@@
A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (Su et al., 2024), enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (Touvron et al., 2023) and Gemma (Team et al., 2024).
@@
  • We introduce the Tensor ProducT ATTenTion Transformer (T6), a new TPA-based model architecture for sequence modeling. On language modeling experiments, T6 consistently improves validation perplexity and downstream evaluation performance with reduced KV cache size.
  • We show TPA integrates seamlessly with RoPE (Su et al., 2024), facilitating easy adoption in popular foundation model architectures such as LLaMA and Gemma.
@@ 2 Background
In this section, we review two classical forms of attention: Scaled Dot-Product Attention, and Multi-Head Attention (MHA) (Vaswani et al., 2017). More types of attention are detailed in the Appendix E, including Multi-Query Attention (MQA) (Shazeer, 2019), and Grouped Query Attention (GQA) (Ainslie et al., 2023), as well as a recent method called Multi-head Latent Attention (MLA) used in DeepSeek-V2 (Liu et al., 2024a) and DeepSeek-V3 (Liu et al., 2024b). We also introduce Rotary Position Embedding (RoPE, Su et al. (2024)), which is commonly used in recent works of large language models.
@@ 2.2 Multi-Head Attention (MHA)
$\mathbf{Q}_{t,i}$ $=(\bm{W}_{i}^{Q})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$
$\mathbf{K}_{t,i}$ $=(\bm{W}_{i}^{K})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$
$\mathbf{V}_{t,i}$ $=(\bm{W}_{i}^{V})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$
@@ 2.3 Rotary Position Embedding (RoPE)
Many recent LLMs use rotary position embedding (RoPE; Su et al., 2024) to encode positional information in the query/key vectors. Specifically, let $\mathrm{RoPE}_{t}$ denote the rotation operator $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ corresponding to the $t$-th position. $\mathbf{T}_{t}$ is a block-diagonal matrix, which consists of block-diagonal matrix$\begin{pmatrix}\cos(t\theta_{j})&-\sin(t\theta_{j})\\ \sin(t\theta_{j})&\cos(t\theta_{j})\end{pmatrix}$, $j\in\{1,\cdots,d_{h}/2\}$, where $\{\theta_{j}\}$ are pre-defined frequency parameters, e.g., $\theta_{j}=1/10000^{2j/d_{h}}$. Then we define
@@ 3.1 Tensor Factorization of Queries, Keys, and Values
$\mathbf{Q}_{t}$ $=\frac{1}{R_{Q}}\sum_{r=1}^{R_{Q}}\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),$ (3.1)
$\mathbf{K}_{t}$ $=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),$ (3.2)
$\mathbf{V}_{t}$ $=\frac{1}{R_{V}}\sum_{r=1}^{R_{V}}\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\;\otimes\;\mathbf{b}^{V}_{r}(\mathbf{x}_{t}),$ (3.3)
where $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h}$, $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}}$. Hence, for queries, each tensor product $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\colon\mathbb{R}^{h}\times\mathbb{R}^{d_{h}}\to\mathbb{R}^{h\times d_{h}}$ adds up to form the query slice $\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}$. Similarly, analogous definitions apply to key slice $\mathbf{K}_{t}$ and value slice $\mathbf{V}_{t}$.
@@
$\mathbf{a}^{Q}(\mathbf{x}_{t})=\bm{W}^{a^{Q}}\,\mathbf{x}_{t}\in\mathbb{R}^{R_{q}\cdot h},~{}\mathbf{b}^{Q}(\mathbf{x}_{t})=\bm{W}^{b^{Q}}\,\mathbf{x}_{t}\in\mathbb{R}^{R_{q}\cdot d_{h}},$
@@ 3.2 RoPE Compatibility and Acceleration
$\widetilde{\mathbf{K}}_{t}$ $=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{(r)}(\mathbf{x}_{t})\otimes\operatorname{RoPE}_{t}\bigl{(}\mathbf{b}^{K}_{(s)}(\mathbf{x}_{t})\bigr{)}$
$=\frac{1}{R_{K}}\mathbf{A}_{K}(\mathbf{x}_{t})^{\top}\operatorname{RoPE}_{t}\bigl{(}\mathbf{B}_{K}(\mathbf{x}_{t})\bigr{)}.$
@@
$\operatorname{RoPE}({\mathbf{Q}}_{t})=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\,\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t}),$ (3.7)
where $\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t})=\operatorname{RoPE}_{t}\bigl{(}\mathbf{B}_{Q}(\mathbf{x}_{t})\bigr{)}$. In addition, assume $\mathbf{Q}_{t}$ and $\mathbf{K}_{s}$ are factorized by TPA and then rotated by $\operatorname{RoPE}_{t},\operatorname{RoPE}_{s}$. Let $\widetilde{\mathbf{Q}}_{t}=\operatorname{RoPE}_{t}(\mathbf{Q}_{t})$ and $\widetilde{\mathbf{K}}_{s}=\operatorname{RoPE}_{s}(\mathbf{K}_{s})$. Then we have
@@
$\widetilde{\mathbf{q}}_{t,i}$ $=\operatorname{RoPE}(\mathbf{q}_{t,i})=\mathbf{T}_{t}\mathbf{q}_{t,i}\in\mathbb{R}^{d_{h}}$
$\quad\widetilde{\mathbf{k}}_{s,i}$ $=\operatorname{RoPE}(\mathbf{k}_{s,i})=\mathbf{T}_{s}\mathbf{k}_{s,i}\in\mathbb{R}^{d_{h}}.$
Theorem 1 indicates that TPA does not break RoPE’s relative translational property. We prove Theorem 1 in Appendix C.1. In short, $\operatorname{RoPE}_{t}$ acts as a block-diagonal orthogonal transform (i.e., a matrix $\mathbf{T}_{t}$) on $\mathbf{B}_{Q}(\mathbf{x}_{t})$. Consequently, $\mathbf{A}_{Q}(\mathbf{x}_{t})$ remains unchanged, while each column of $\mathbf{B}_{Q}(\mathbf{x}_{t})$ is rotated appropriately, preserving the TPA structure.